# CURVALID: Geometrically-guided Adversarial Prompt Detection

### v1

@@ Abstract
Adversarial prompts capable of jailbreaking large language models (LLMs) and inducing undesirable behaviours pose a significant obstacle to their safe deployment. Current mitigation strategies rely on activating built-in defence mechanisms or fine-tuning the LLMs, but the fundamental distinctions between adversarial and benign prompts are yet to be understood. In this work, we introduce CurvaLID, a novel defense framework that efficiently detects adversarial prompts by leveraging their geometric properties. It is agnostic to the type of LLM, offering a unified detection framework across diverse adversarial prompts and LLM architectures. CurvaLID builds on the geometric analysis of text prompts to uncover their underlying differences. We theoretically extend the concept of curvature via the Whewell equation into an $n$-dimensional word embedding space, enabling us to quantify local geometric properties, including semantic shifts and curvature in the underlying manifolds. Additionally, we employ Local Intrinsic Dimensionality (LID) to capture geometric features of text prompts within adversarial subspaces. Our findings reveal that adversarial prompts differ fundamentally from benign prompts in terms of their geometric characteristics. Our results demonstrate that CurvaLID delivers superior detection and rejection of adversarial queries, paving the way for safer LLM deployment. The source code can be found at https://github.com/Cancanxxx/CurvaLID.
Curvature, Local intrinsic dimension, Large Language Model, Adversarial Defense
@@ 1 Introduction
Large Language Models (LLMs) are widely used in real-world applications such as education, customer service, and legal analysis (Yao et al., 2024). However, adversarial prompts exploit vulnerabilities of LLMs to produce unintended and harmful responses (Wallace et al., 2019; Shen et al., 2023; Deng et al., 2023a). Therefore, ensuring their safety against adversarial prompts is essential to prevent harmful outputs, such as biased language, misinformation, or content inciting physical harassment.
@@
Current defences against adversarial prompts mainly rely on prompt engineering or adversarial training. Input perturbation techniques, like IntentionAnalysis (Zhang et al., 2024a) and SmoothLLM (Robey et al., 2023), operate by modifying the prompt and examining whether the altered version can successfully trigger the LLM’s built-in defence mechanisms. The effectiveness of these methods depends on the degree of perturbation and the robustness of the LLM’s built-in defense mechanisms. Meanwhile, adversarial training approaches, which involve fine-tuning the model to resist adversarial inputs, often struggle to scale to larger LLMs. For example, Latent Adversarial Training (LAT) cannot be easily applied to LLMs exceeding 10 billion parameters (Sheshadri et al., 2024). Given that popular LLMs, such as GPT-3 and PaLM 2, have 175 billion and 340 billion parameters respectively, developing scalable defenses for these large models is essential to ensure their reliability and safety (Anil et al., 2023; Brown, 2020). These models are widely used in applications like automated content generation, customer support, and medical diagnostics, where ensuring robustness against adversarial attacks is critical.
Additionally, these existing solutions are inherently tied to the internal architecture of the targeted LLM, limiting their generality. They do not guarantee consistent performance across different adversarial prompts and varying models. Most importantly, the fundamental differences between adversarial and benign prompts remain poorly understood, underscoring the need for a more generalizable and theoretically grounded defence strategy. Understanding these differences is critical, as adversarial prompts exploit specific vulnerabilities that vary across LLMs, making it difficult for current defenses to address them comprehensively. A deeper understanding could enable the development of effective, robust strategies across diverse models and attack scenarios.
In this work, we introduce CurvaLID, a model-agnostic framework designed to address the limitations of current solutions. CurvaLID provides a defense mechanism that operates independently of the internal architecture of targeted LLMs, ensuring its generality across diverse models. It ensures LLMs safety by detecting and filtering out adversarial prompts before they are being used. CurvaLID distinguishes between benign and adversarial prompts by examining their geometric distinctions. Specifically, it characterizes their Local Intrinsic Dimensionality (LID) at the sentence level and introduces a curvature measure for text at the word level, enabling a comprehensive analysis of the geometric properties of text prompts. LID has previously been shown to characterize adversarial subspaces in the vision domain (Ma et al., 2018), and token-level LID estimation has been explored to measure the truthfulness of LLM-generated text (Yin et al., 2024). However, we find that token-level estimation is insufficient for detecting adversarial prompts, as the local neighbourhood often consists of stop words, providing limited semantic information. To address this, we propose measuring LID at the prompt level—referred to as PromptLID—rather than relying on token-level analysis.
@@
Additionally, we develop TextCurv, a theoretical framework for defining curvature in an $n$-dimensional Euclidean space of word embeddings. By extending the curvature concept from the Whewell equation (Whewell, 1849), we prove that the angle between two tangent vectors is equivalent to the difference in their tangential angles. This provides the foundation for analyzing word-level geometry, enabling us to quantify curvature in text embeddings and capture semantic shifts. Through extensive evaluations, we demonstrate that PromptLID effectively quantifies the high-dimensional local subspaces where adversarial prompts reside, while TextCurv captures curvature at the word level. Together, they reveal semantic shifts and localized structural deviations in the text manifold, providing new insights into the fundamental differences between adversarial and benign prompts.
@@
  • We propose CurvaLID, a model-agnostic detection framework that leverages geometric distinctions between adversarial and benign prompts. By integrating TextCurv and PromptLID, CurvaLID efficiently and effectively detects adversarial prompts, ensuring robustness across various LLMs.
  • We provide theoretical insights to the design of two novel geometric measures leveraged by CurvaLID. TextCurv extends the Whewell equation to $n$-dimensional Euclidean space, enabling the quantification of semantic shifts through word-level curvature. PromptLID analyzes the local intrinsic dimensionality across entire prompts, capturing adversarial subspaces more effectively.
  • CurvaLID successfully detects about 99% of adversarial prompts, outperforming the state-of-the-art defenses on various adversarial prompts by large margins. It is also highly time-efficient, requiring only 0.25 GPU hours of training, whereas existing adversarial training methods have significantly higher computational costs, usually exceeding 100 GPU hours.
@@ 2 Related Work
This section reviews prior research on adversarial attacks and defense mechanisms for LLMs, highlighting their objectives, underlying logic, and key characteristics.
Adversarial attacks on LLM. Adversarial attacks on LLMs involve crafted inputs designed to manipulate models into generating harmful content, like offensive language or dangerous instructions (Zou et al., 2023). These attacks range from a single input (zero-shot) to more complex, continuous dialogue scenarios (multi-shot) (Shen et al., 2023; Dong et al., 2023; Wang et al., 2023). This research focuses on zero-shot text prompt attacks, including techniques like text perturbation that adds gibberish or subtly alters input wording and social-engineered prompts that trick LLMs into harmful behavior (Zou et al., 2023; Schwinn et al., 2023; Chu et al., 2024).
Adversarial defenses for LLM. There are three primary defenses against adversarial attacks on LLMs: input preprocessing, prompt engineering, and adversarial training. Input preprocessing perturbs inputs to disrupt adversarial prompts but may also affect benign ones, with its effectiveness varying depending on the extent of the perturbation and the targeted LLM (Cao et al., 2023; Robey et al., 2023; Yung et al., 2024). Prompt engineering enhances the LLM’s self-defensive capabilities by adding prompts that help expose harmful intent, though its effectiveness varies with different models (Zhang et al., 2024a; Phute et al., 2023; Zhang et al., 2024b). Finally, adversarial training strengthens the LLM’s ability to reject harmful prompts by exposing it to adversarial scenarios during training (Xu et al., 2024b; Jain et al., 2023). However, this approach requires fine-tuning the LLM, making its success dependent on the specific model being protected, with most adversarial training methods confined to white-box models. Additionally, these methods often demand significant computational resources, with training times reaching up to 128 GPU hours (Mazeika et al., 2024) or 12 GPU hours (Sheshadri et al., 2024), and generally exhibit varying effectiveness across different adversarial prompts and LLMs (Sheshadri et al., 2024; Ziegler et al., 2022; Ganguli et al., 2022). Note that our paper belongs to the field of adversarial prompt detection. Unlike existing methods such as perplexity filtering, which rely on LLMs for next-token probability, our method operates independently of the type of the LLM (Hu et al., 2023).
@@ 3 Background and Terminology
This section provides a brief overview of the mathematical definitions of LID and curvature, as well as how they are estimated and utilized in practice.
@@ 3.1 Local Intrinsic Dimension
Local Intrinsic Dimensionality (LID) measures the intrinsic dimensionality of the local neighborhood around a reference sample (Houle, 2017). Compared to Global Intrinsic Dimension (GID) (Tulchinskii et al., 2024; Pope et al., 2021a), which measures the degree of the $d$-dimension of the global manifold of a data subset, LID focuses on the local neighborhood of given points. Thus, LID is particularly useful in analyzing high-dimensional data with varying dimensionality across the dataset.
@@
$\operatorname{LID}^{*}_{F}$ is the theoretical definition, and in practice, this quantity needs to be estimated (Levina & Bickel, 2004; Tempczyk et al., 2022). In machine learning applications, estimation of LID requires a distance measure and a set of reference points to select nearest neighbors (NN). Following prior work (Gong et al., 2019; Ansuini et al., 2019; Pope et al., 2021b; Zhou et al., 2024a; Huang et al., 2024, 2025), we use the Euclidean distance. The representation of a data point, along with the chosen reference points, significantly influences how LID is interpreted. In adversarial prompt detection, the representation and neighborhood definition directly affect the ability to distinguish between clean and adversarial prompts. Among existing estimators, we use the Method of Moments (MoM) (Amsaleg et al., 2015) due to its simplicity.
@@ 3.2 Curvature
The intuition of curvature is how quickly a curve changes direction. In geometry, we can visualize curvature through an osculating circle. Curvature can be measured at a given point by fitting a circle to the curve on which the point resides (Kline, 1998). The formal definition is as follows:
@@
[Curvature measured by osculating circle](Kline, 1998)
The osculating circle for some point $P$ on a curve is defined as the circle tangent to $P$ and passing each other point on the curve. We denote the radius of the osculating circle as $R$. Thus, the curvature $\kappa$ is defined as:
@@
[Curvature in differential geometry](Shifrin, 2015) Suppose the curve $\alpha$ is parametrized by arc length $s$ and $\mathbf{T}(s)$ is the unit tangent vector to the curve. We define curvature as
@@
Curvature is also defined and utilized in physics. The Frenet-Serret formulas relate curvature to torsion, tangent, normal, and binormal unit vectors (Frenet, 1852). In the Frenet-Serret formulas, the curvature describes the rotational speed along a curve, which is relevant in kinetics and trajectory applications (Huang et al., 2023). It is also used in autonomous driving, robotics, and quantum computing (Hallgarten et al., 2024; Alsing & Cafaro, 2023; Shabana, 2023).
4 Geometric Analysis and CurvaLID Framework
We aim to develop geometric measures that can effectively characterize benign and adversarial prompts at both the prompt and word levels, and utilize these measurements for adversarial prompt detection. We specifically define our problem statement as follows:
Let $\mathcal{D}=\{(x^{i},y^{i})\}_{i=1}^{n}$ be a labeled dataset comprising $n$ i.i.d. samples $x^{i}$, where each sample is associated with a label $y^{i}$. In this context, each input $x^{i}$ represents a text prompt, with the corresponding label $y^{i}\in\{0,1\}$ indicating whether the prompt is benign ($y^{i}=0$), or adversarial ($y^{i}=1$). Let $\mathcal{M}(x)$ denote the geometric measure applied to a prompt $x$, where $\mathcal{M}(x)$ is composed of two complementary components: the prompt-level measure $\text{PromptLID}(x)$ and the word-level measure $\text{TextCurv}(x)$. These measures are then utilized within an adversarial prompt detection algorithm, formalized as a classification problem where the objective is to minimize the empirical error between the ground-truth labels and the predictions:
@@ 4 Geometric Analysis and CurvaLID Framework
where $\ell(\cdot)$ denotes the cross-entropy loss function, and $h(\mathcal{M}(x))$ is the classifier applied to the geometric measures $\mathcal{M}(x)=(\text{PromptLID}(x),\text{TextCurv}(x))$. Next, we formally define PromptLID and TextCurv, detailing how they explore the geometric properties at the prompt and word levels, respectively. Finally, we provide an overview of CurvaLID, our adversarial prompt detection model.
@@ 4.1 PromptLID: LID estimation at the prompt-level
To characterize the prompt-level geometric properties of benign and adversarial prompts, we propose PromptLID, an LID estimation based on prompt representations obtained from a trained CNN. We first train a model $g$ (CNN) to perform a $k$-class classification task, where the goal is to determine which benign dataset a given prompt belongs to. This involves learning a function $g:\mathcal{B}\rightarrow\mathcal{Q}$ to map the input space $\mathcal{B}$ to the label space $\mathcal{Q}$. The label space is defined as $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{k}\}$, where $k$ is the number of types of benign datasets and is equal to the cardinality of $\mathcal{Q}$. Given a benign prompt dataset =$\mathcal{B}=\{(b,q)^{i}\}_{i=1}^{n}$, where $b$ is the benign prompt and $q$ is its corresponding label, the model learns to classify each prompt into its respective dataset. The objective function used is categorical cross-entropy, as the task involves multi-class classification. The representation $\bm{z}_{1}$, derived from the penultimate dense layer, encodes the prompt as a single vector, which is then used to calculate the PromptLID. The PromptLID expands on the MoM estimation (Amsaleg et al., 2015) of LID on the prompts’ representation in $\bm{z}_{1}$.
Definition 4.1 (PromptLID).
The PromptLID of a given prompt $\bm{x}$ is defined as:
$PromptLID=-k\cdot\frac{\mu_{k}}{\mu_{k}-w^{k}},$
where $k$ is a locality hyperparameter representing the number of nearest neighbors, $\mu_{k}$ is the averaged distance from the prompt representation $\bm{z}_{1}$ of $\bm{x}$ to its $k$-nearest neighbors, and $w^{k}$ is the distance to the $k$-th nearest neighbor.
Given that adversarial prompts often manipulate the high-dimensional space of word embeddings to target rarely encountered subspaces (Szegedy, 2013; Ma et al., 2018), they exhibit distinct geometric characteristics. Adversarial prompts are expected to exhibit higher PromptLID as they push the embeddings into regions of the feature space that are less well-defined and more complex than typical benign inputs. As shown in Section 5.2, PromptLID effectively captures this behavior, highlighting its ability to distinguish between benign and adversarial prompts.

### v2

# Geometry-Guided Adversarial Prompt Detection via Curvature and Local Intrinsic Dimension

@@ Abstract
Adversarial prompts are capable of jailbreaking frontier large language models (LLMs) and inducing undesirable behaviours, posing a significant obstacle to their safe deployment. Current mitigation strategies primarily rely on activating built-in defence mechanisms or fine-tuning LLMs, both of which are computationally expensive and can sacrifice model utility. In contrast, detection-based approaches are more efficient and practical for deployment in real-world applications. However, the fundamental distinctions between adversarial and benign prompts remain poorly understood. In this work, we introduce CurvaLID, a novel defence framework that efficiently detects adversarial prompts by leveraging their geometric properties. It is agnostic to the type of LLM, offering a unified detection framework across diverse adversarial prompts and LLM architectures. CurvaLID builds on the geometric analysis of text prompts to uncover their underlying differences. We theoretically extend the concept of curvature via the Whewell equation into an $n$-dimensional word embedding space, enabling us to quantify local geometric properties, including semantic shifts and curvature in the underlying manifolds. To further enhance our solution, we leverage Local Intrinsic Dimensionality (LID) to capture complementary geometric features of text prompts within adversarial subspaces. Our findings show that adversarial prompts exhibit distinct geometric signatures from benign prompts, enabling CurvaLID to achieve near-perfect classification and outperform state-of-the-art detectors in adversarial prompt detection. CurvaLID provides a reliable and efficient safeguard against malicious queries as a model-agnostic method that generalises across multiple LLMs and attack families.
@@ 1 Introduction
Frontier Large Language Models (LLMs) are widely used in real-world applications such as education, finance, and legal analysis (Yao et al., 2024). However, adversarial prompts exploit vulnerabilities of LLMs to produce unintended and harmful responses (Wallace et al., 2019; Shen et al., 2023; Deng et al., 2023a). Therefore, ensuring their safety against adversarial prompts is essential to prevent harmful outputs, such as bias, misinformation, or content inciting physical harassment.
@@
Current defences against adversarial prompts rely on prompt engineering or adversarial training. Input perturbation techniques, like Intentionanalysis (Zhang et al., 2024a) and SmoothLLM (Robey et al., 2023), modify the prompt and examine whether the altered version can successfully trigger the LLM’s built-in safety mechanisms. The effectiveness of these methods depends on the degree of perturbation and robustness of the LLM’s safety alignment. Meanwhile, adversarial training approaches, which fine-tune models to resist adversarial inputs, often struggle to scale to larger LLMs. For example, Latent Adversarial Training (LAT) cannot be easily applied to LLMs exceeding 10 billion parameters (Sheshadri et al., 2024). Given that popular LLMs, such as GPT-3 and PaLM 2, have 175 billion and 340 billion parameters respectively, developing scalable defences for these large models is essential to ensure their reliability and safety (Anil et al., 2023; Brown, 2020).
Existing solutions are inherently tied to the internal architecture and safety alignment training of the targeted LLM, limiting their generality. They do not guarantee consistent performance across different adversarial prompts and varying models (Chao et al., 2023; Shen et al., 2023; Zhou et al., 2024b). In parallel, Llama Guard (Inan et al., 2023) and the constitutional classifier (Sharma et al., 2025) are trained to detect harmful adversarial prompts and block them from reaching the LLM. However, both approaches rely on human annotations to differentiate harmful from benign inputs. More importantly, the underlying distinctions between adversarial and benign prompts remain insufficiently understood, underscoring the need for defence strategies that are both generalizable and theoretically grounded.
In this work, we introduce CurvaLID, an LLM-agnostic framework designed to explore generalizable solutions by uncovering the fundamental geometric differences between adversarial and benign prompts. CurvaLID provides a defence mechanism that operates independently of the internal architecture of LLMs, ensuring its generality across diverse models. It enhances LLM safety by preemptively rejecting adversarial inputs.
First, we introduce PromptLID, a sentence-level Local Intrinsic Dimensionality (LID) measure that effectively captures the geometric properties of text prompts. PromptLID calculates LID using a sentence-level-defined local neighbourhood, unlike traditional approaches (Ma et al., 2018; Yin et al., 2024) that rely on token-level neighbourhoods, which are easily influenced by local noise and often dominated by stop words that provide limited semantic information. Prior work has applied LID to characterise adversarial subspaces in the vision domain (Ma et al., 2018) and to assess the truthfulness of LLM outputs using token-level estimation (Yin et al., 2024), but these methods are not agnostic to specific models or LLM architectures, limiting their generalizability. PromptLID addresses these issues by offering a sentence-level, model-agnostic characterization of prompt-level geometry, enabling more robust adversarial prompt detection.
@@
Second, we develop TextCurv, a theoretical framework for defining curvature in an $n$-dimensional Euclidean space of word embeddings. By extending the curvature concept from the Whewell equation (Whewell, 1849), we prove that the angle between two tangent vectors is equivalent to the difference in their tangential angles. This provides the foundation for analyzing word-level geometry, enabling us to quantify curvature in text embeddings and capture semantic shifts.
Through extensive evaluations, we demonstrate that PromptLID effectively quantifies the high-dimensional local subspaces where adversarial prompts reside, while TextCurv captures curvature at the word level. Together, they complement each other in revealing semantic shifts and localized structural deviations in the text manifold, providing new insights into the fundamental differences between adversarial and benign prompts.
@@
  • We propose CurvaLID, a LLM-agnostic detection framework that leverages geometric distinctions between adversarial and benign prompts. By integrating TextCurv and PromptLID, CurvaLID efficiently and effectively detects adversarial prompts, ensuring safety across various LLMs.
  • We provide theoretical insights into the design of two novel geometric measures leveraged by CurvaLID. TextCurv extends the Whewell equation to $n$-dimensional Euclidean space, enabling quantification of semantic shifts through word-level curvature. PromptLID analyses the local intrinsic dimensionality across entire prompts, capturing adversarial subspaces effectively.
  • CurvaLID successfully detects about 99% of adversarial prompts, outperforming state-of-the-art defences by over 10% in attack success rate reduction across multiple LLMs and adversarial attacks. It is highly time-efficient, requiring only 0.25 GPU hours of training, whereas existing adversarial training methods have significantly higher computational costs, usually exceeding 100 GPU hours.
@@ 2 Related Work
This section reviews prior research on adversarial attacks and defence mechanisms for LLMs, highlighting their objectives, underlying logic, and key characteristics.
Adversarial attacks on LLM. Adversarial attacks on LLMs involve crafted inputs designed to manipulate models into generating harmful content, like offensive language or dangerous instructions (Zou et al., 2023). These attacks range from a single input (zero-shot) to more complex, continuous dialogue scenarios (multi-shot) (Shen et al., 2023; Dong et al., 2023; Wang et al., 2023). This research focuses on zero-shot text prompt attacks, including techniques like text perturbation that adds gibberish or subtly alters input wording and social-engineered prompts that trick LLMs into harmful behaviour (Zou et al., 2023; Schwinn et al., 2023; Chu et al., 2024).
Adversarial defences for LLM. There are three primary defences against adversarial attacks on LLMs: input preprocessing, prompt engineering, and adversarial training. Input preprocessing perturbs inputs to disrupt adversarial prompts but may also affect benign ones, with effectiveness depending on perturbation level and targeted LLM (Cao et al., 2023; Robey et al., 2023; Yung et al., 2024). Prompt engineering augments self-defensive behaviour by adding prompts to expose harmful intent, though performance varies across models (Phute et al., 2023; Zhang et al., 2024a; b). Finally, adversarial training strengthens the LLM’s ability to reject harmful prompts by exposing models to adversarial cases during training (Jain et al., 2023; Xu et al., 2024b). However, this approach requires fine-tuning the LLM, making its success dependent on the specific model being protected, with most adversarial training methods confined to white-box models. Additionally, these methods often demand significant computational resources, with training times reaching up to 128 GPU hours (Mazeika et al., 2024) or 12 GPU hours (Sheshadri et al., 2024), and generally exhibit varying effectiveness across different adversarial prompts and LLMs (Ziegler et al., 2022; Ganguli et al., 2022; Sheshadri et al., 2024). Note that our paper belongs to the field of adversarial prompt detection. Unlike existing methods such as perplexity filtering, which rely on LLMs for next-token probability, our method operates independently of the type of the LLM (Hu et al., 2023).
@@ 3 Background and Terminology
This section provides a brief overview of the mathematical definitions of LID and curvature.
@@ 3.1 Local Intrinsic Dimension
Local Intrinsic Dimensionality (LID) measures the intrinsic dimensionality of the local neighbourhood around a reference sample (Houle, 2017). Compared to Global Intrinsic Dimension (GID) (Tulchinskii et al., 2024; Pope et al., 2021b), which measures the degree of the $d$-dimension of the global manifold of a data subset, LID focuses on the local neighbourhood of given points. Thus, LID is particularly useful in analyzing high-dimensional data with varying dimensionalities across the dataset.
@@
$\operatorname{LID}^{*}_{F}$ is the theoretical definition, and in practice, has to be estimated (Levina & Bickel, 2004; Tempczyk et al., 2022). Estimation of LID requires a distance measure and a set of reference points to select nearest neighbors (NN). Following prior work (Gong et al., 2019; Ansuini et al., 2019; Pope et al., 2021a; Zhou et al., 2024a; Huang et al., 2024; 2025), we use Euclidean distance. The representation of a data point, along with the chosen reference points, significantly influences how LID is interpreted. In adversarial prompt detection, the representation and neighbourhood definition directly affect the ability to distinguish between clean and adversarial prompts. Among existing estimators, we use the Method of Moments (MoM) (Amsaleg et al., 2015) for its simplicity.
@@ 3.2 Curvature
The intuition of curvature is how quickly a curve changes direction. In geometry, we can visualise curvature through an osculating circle. Curvature can be measured at a given point by fitting a circle to the curve on which the point resides (Kline, 1998). The formal definition is as follows:
@@
[Curvature measured by osculating circle] (Kline, 1998) The osculating circle at a point $P$ on a curve is the circle tangent at $P$ and passing through nearby points on the curve. Let $R$ be its radius. The curvature $\kappa$ is then defined as:
@@
[Curvature in differential geometry](Shifrin, 2015) Suppose curve $\alpha$ is parametrized by arc length $s$ and $\mathbf{T}(s)$ is the unit tangent vector to the curve. We define curvature as
@@
Curvature is also defined and utilized in physics. The Frenet-Serret formulas relate curvature to torsion, tangent, normal, and binormal unit vectors (Frenet, 1852). In the Frenet-Serret formulas, the curvature describes the rotational speed along a curve, which is relevant in kinetics and trajectory applications (Huang et al., 2023). It is also used in autonomous driving, robotics, and quantum computing (Alsing & Cafaro, 2023; Shabana, 2023; Hallgarten et al., 2024).
4 Geometric analysis and CurvaLID
We aim to develop geometric measures that effectively characterise both benign and adversarial prompts at the prompt and word levels. These measures are then utilized for adversarial prompt detection, formulated as a binary classification task defined as follows.
Let $\mathcal{D}=\{(x^{i},y^{i})\}_{i=1}^{n}$ be a labelled dataset comprising $n$ i.i.d. samples $x^{i}$, where each sample is associated with a label $y^{i}$. In this context, each input $x^{i}$ represents a text prompt, with the corresponding label $y^{i}\in\{0,1\}$ indicating whether the prompt is benign ($y^{i}=0$), or adversarial ($y^{i}=1$). Let $\mathcal{M}(x)$ denote the geometric measure applied to a prompt $x$, where $\mathcal{M}(x)$ is composed of two complementary components: the prompt-level measure $\text{PromptLID}(x)$ and the word-level measure $\text{TextCurv}(x)$. These measures are then utilized within an adversarial prompt detection algorithm, formalized as a classification problem where the objective is to minimize the empirical error between the ground-truth labels and the predictions:
@@ 4 Geometric analysis and CurvaLID
where $\ell(\cdot)$ denotes the cross-entropy loss function, and $h(\mathcal{M}(x))$ is the classifier applied to the geometric measures $\mathcal{M}(x)=(\text{PromptLID}(x),\text{TextCurv}(x))$. Alternatively, the defender may use classical outlier detection methods without access to adversarial prompts during training. In both cases, the detector operates on the same geometric measures. Next, we formally define PromptLID and TextCurv, detailing how they explore geometric properties at the prompt and word levels, respectively. Finally, we provide an overview of CurvaLID, our adversarial prompt detection model.
@@ 4.1 PromptLID: LID estimation at the prompt-level
To characterise the prompt-level geometric properties of benign and adversarial prompts, we propose PromptLID, an LID estimation based on prompt representations obtained from a trained CNN. We first train a model $g$ (CNN) to perform a $k$-class classification task, where the goal is to determine which benign dataset a given prompt belongs to. This involves learning a function $g:\mathcal{B}\rightarrow\mathcal{Q}$ to map the input space $\mathcal{B}$ to the label space $\mathcal{Q}$. The label space is defined as $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{k}\}$, where $k$ is the number of types of benign datasets and is equal to the cardinality of $\mathcal{Q}$. Given a benign prompt dataset $\mathcal{B}=\{(b,q)^{i}\}_{i=1}^{n}$, where $b$ is the benign prompt and $q$ is its corresponding label, the model learns to classify each prompt into its respective dataset. The objective function used is categorical cross-entropy, as the task involves multi-class classification. The representation $\bm{z}_{1}$, derived from the penultimate dense layer, encodes the prompt as a single vector, which is then used to calculate the PromptLID. The PromptLID expands on the MoM estimation (Amsaleg et al., 2015) of LID on the prompts’ representation in $\bm{z}_{1}$.
Definition 4.1.
[PromptLID] The PromptLID of a prompt $\bm{x}$ is defined as:
$\mathrm{PromptLID}=-k\cdot\frac{\mu_{k}}{\mu_{k}-w^{k}},$
where $k$ is the number of nearest neighbors, $\mu_{k}$ is the mean distance from the prompt representation $\bm{z}_{1}$ to its $k$-nearest neighbors, and $w^{k}$ is the distance to the $k$-th neighbor.
Given that adversarial prompts often manipulate the high-dimensional space of word embeddings to target rarely encountered subspaces (Szegedy, 2013; Ma et al., 2018), they exhibit distinct geometric characteristics. Adversarial prompts are expected to exhibit higher PromptLID as they push the embeddings into regions of the feature space that are less well-defined and more complex than typical benign inputs. As shown in Section 5.2, PromptLID effectively captures this behaviour, highlighting its ability to distinguish between benign and adversarial prompts.